OpenStudy (anonymous):

verify: tan^2 θ+1 / 1-sin^2 θ =sec^4θ

OpenStudy (princeharryyy):

tan^2theta + 1 = sec^2theta and 1- sin^2theta = cos^2theta so, sec^2theta/cos^2theta = sec^4theta, as cos^2theta = 1/sec^2theta

OpenStudy (princeharryyy):

proved @lajessica.

OpenStudy (anonymous):

thank you so much !

OpenStudy (princeharryyy):

:)

OpenStudy (princeharryyy):

yw